My question was about arithmetic series. Sn = -126, t1= -1, and tn = -20
How do I determine n?

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

My question was about arithmetic series. Sn = -126, t1= -1, and tn = -20
How do I determine n?
-------
Using S(n) = -126 you get: -126 = (n/2)(-1-20)
-----
-252 = -21n
n = 12
===========
Cheers
Stan H.
===========

Answer by mananth(16946)

Sn = -126, t1= -1, and tn = -20
How do I determine n?
tn = a+(n-1)d
-20 =-1+(n-1)d
-19=(n-1)d
Sn= n/2[2a+(n-1)d]
-126=n/2[2(-1)+(n-1)d]
substitute (n-1)d =-19
-252=n[-2-19)]
-252=n[-21]
-252/-21=n
n=22